# Are Language Models Actually Useful for Time Series Forecasting?

### v1

@@ Abstract
Large language models (LLMs) are being applied to time series tasks, particularly time series forecasting. However, are language models actually useful for time series? After a series of ablation studies on three recent and popular LLM-based time series forecasting methods, we find that removing the LLM component or replacing it with a basic attention layer does not degrade the forecasting results—in most cases the results even improved. We also find that despite their significant computational cost, pretrained LLMs do no better than models trained from scratch, do not represent the sequential dependencies in time series, and do not assist in few-shot settings. Additionally, we explore time series encoders and reveal that patching and attention structures perform similarly to state-of-the-art LLM-based forecasters. [^1] All code, environments, and data to reproduce our work are available in this anonymized repository: https://github.com/BennyTMT/TS_Models
@@ 1 Introduction
Time series analysis is a critical problem across many domains, including disease propagation forecasting [7], retail sales analysis [3], healthcare [23, 15] and finance [28]. A great deal of recent work in time series analysis (constituting repositories with more than 1200 total stars on GitHub) has focused on adapting pretrained large language models (LLMs) to classify, forecast, and detect anomalies in time series [13, 42, 19, 4, 5, 29, 12, 37, 14]. These papers posit that language models, being advanced models for sequential dependencies in text, may generalize to the sequential dependencies in time series data. This hypothesis is unsurprising given the popularity of language models in machine learning research writ large. So to what extent are language models really beneficial for traditional time series tasks?
Our main claim is simple but profound: popular methods for adapting language models for time series forecasting perform the same or worse than basic ablations, yet require orders of magnitude more compute. Derived from extensive ablations, these findings reveal a worrying trend in contemporary time series forecasting literature. Our goal is not to imply that language models will never be useful for time series. In fact, recent works point to many exciting and promising ways that language and time series interact, like time series reasoning [22] and social understanding [6]. Rather, we aim to highlight surprising findings that existing methods do very little to use the innate reasoning power of pretrained language models on established time series tasks.
@@
  • We study the impact of an LLM’s pretraining by re-initializing their weights prior to forecasting. We find that this has no impact on forecasting performance. Additionally, in shuffling input time series, we find no evidence the LLMs successfully transfer sequence modeling abilities from text to time series and no indication that they help in few-shot settings.
@@ 2 Related Work
Here, we summarize the key related works for our paper. They can be broadly classified into three sections: (i) time series forecasting using LLMs; (ii) encoders in LLM time series models; and (iii) smaller and efficient neural models for time-series.
Time Series Forecasting Using LLMs. Recently, with the development of Large Language Models (LLMs) [9, 26, 31] and their demonstrated multi-modal capabilities, more researchers have successfully applied LLMs to time series forecasting tasks [12, 14, 5, 4]. Chang et al.,  [5] used finetuning the transformer module and positional encoding in GPT-2 to align pre-trained LLMs with time series data for forecasting tasks. Zhou et al. [42] proposed a similar finetuning method, named “OneFitAll”, for time series forecasting with GPT-2. Additionally, Jin et al. [13] introduced a reprogramming method to align LLM’s Word Embedding with time series embeddings, showing good representation of time series data on LLaMA [31]. Similarly, LlaTA [19] and Test [29] adapted word embeddings to enable LLMs to forecast time series data effectively. In addition to time-series forecasting models, Liu et al. [20] show that these models can be extended to classifying health-time series, such as heart-rate and daily-footsteps. These models have also been shown to outperform supervised neural models in few-shot settings.
Encoders in LLM Time Series Models. In order for an LLM to learn from text it must first be discretized and encoded as word tokens which are $1\times d$ vectors [9, 26, 31]. Similarly, LLM-based methods for time series learn discrete time series tokens. One method this is to segment the time series into overlapping patches, which effectively shortens the time series while retaining its features [13, 42, 5, 4, 25, 24]. Another method involves decomposing a method based on its trend, seasonal components, and residual components [4, 25]. Lastly, Liu et al. [19] feed the multivariate time series using a Transformer to enable different channels to learn the dynamics of other channels. These embedding procedures are followed by a linear neural network layer that projects the time series encoding to the same dimensions used by the pre-trained LLM.
Smaller and Efficient Neural Models. In addition to LLMs, there has been a large body of research on smaller yet efficient frameworks that outperform their bulky counterparts in time series forecasting [17, 39, 30, 21, 2]. For example, Zeng et al. [38] present DLinear, an incredibly simple model that combines decomposition techniques and achieves better forecasting performance than state-of-the-art transformer-based time series architectures at the time, such as Informer [40], FEDformer [41], and Autoformer [35]. Furthermore, Xu et al. [36] introduces a lightweight model with only 10k parameters, which captures both amplitude and phase information in the time-series to outperform transformer-based models.
@@ 3 Experimental Setup
LLaTA [19] GPT-2 LoRA Fine-Tune
@@ 3.1 Reference Methods for Language Models and Time Series
  • LLaTA [19]: LLaTA embeds the input time series by treating each channel as a token. One half of the architecture is a “textual branch” which uses cross attention to align the time series representation with a low dimensional representation of the language model’s word embeddings. This representation is then passed through a pretrained, frozen language model to obtain a “textual prediction”. Simultaneously, a “temporal” branch learns a low-rank adapter for a pretrained language model based on the input time series to produce a “temporal prediction” which is used for inference. The model includes additional loss terms that enforce similarity between these representations.
@@ 4 Results
LLaTA [19] Model $\rightarrow$ LLaTA w/o LLM LLM2Attn LLM2Trsf From Original Paper
@@
Table 3: Forecasting performance of all models – Time-LLM, LLaTA, and OneFitsAll and results from our ablations. All results are averaged across different prediction lengths, though full results are available in Appendix E.1. Results in Red denote the best-performing model. # Wins refers to the number of times the method performed best, and # Params is the number of model parameters. “-” means the dataset is not included in the original paper.
@@ 4.1 Do pretrained language models contribute to forecasting performance? (RQ1)
Our results show that pretrained LLMs are not useful for time series forecasting tasks yet. Overall, as shown in Table 3, across 8 datasets and two metrics, ablations out perform Time-LLM methods in $/26$ 26 cases, LLaTA in $22/26$ cases, and OneFitsAll in $19/26$ cases. We averaged results over different predicting lengths, as in [42, 13, 19]. Across all prediction lengths (eight datasets and four prediction lengths) ablations outperformed Time-LLM, LLaTA, and OneFitsAll in $35/40$, $31/40$, and $29/40$ cases as measured by MAE, respectively. To ensure a fair comparison, we also report results from each method’s original paper alongside our replication. For specific results refer to Appendix E.1. To better evaluate the effectiveness of LLMs and ablation methods, we include 95% bootstrapped confidence intervals for each task. In tasks where LLMs performed better, such as OneFitsAll with ETTh1, shown in  Figure 2, there is still substantial overlap in the confidence intervals with the ablation method “w/o LLM” in MAE. Other datasets results and MSE metrics are shown in Figure 5 and Figure 6 in the Appendix. To summarize, our results on the evaluation above, it is hard to conclude that LLMs are effective in time series forecasting.
Figure 2: In the above examples, only OneFitsAll “w/ LLM” performs better than the ablation methods on ETTh1, but there is substantial overlap in bootstraped confidence intervals. The figures show the comparison of OneFitsAll, LLaTA, and Time-LLM using LLMs and ablations (i.e., w/o LLM, LLM2Attn, and LLM2Trsf) on ETTh1, ETTm2, and Electricity, and the vertical dashed lines represent the results from the original work. Others Figures for MSE and other datasets are available in Figure 5 and Figure 6 in the Appendix.
Figure 3: Ablation methods consume less time for inference while providing better forecasting performance. The figure above shows the inference time and prediction accuracy of Time-LLM, OneFitsAll, and LLaTA on ETTm2, Traffic, and Electricity datasets, averaged across prediction lengths. For more datasets and MSE metrics refer to Figure 7 and Figure 8 in the Appendix.
Method Time-LLM (LLaMA) OneFitsAll (GPT-2) LLaTA (GPT-2)
@@ 4.2 Are LLM-based methods worth the computational cost? (RQ2)
In the previous section, we showed that LLMs do not meaningfully improve performance on time series forecasting tasks. Here, we evaluate the computational intensity of these methods with their nominal performance in mind. The language models in our reference methods use hundreds of millions and sometimes billions of parameters to perform time series forecasting. Even when the parameters of the language models are frozen they still contribute to substantial overhead during training and inference. For instance, Time-LLM has 6642 M parameters and takes 3003 minutes to train on the Weather dataset whereas ablation methods have only 0.245 M parameters and take 2.17 minutes on average. Information about training other methods on ETTh1 and Weather datasets are shown in Table 4. In the case of inference time, we divide by the maximum batch size to give an estimate of inference time per example. Time-LLM, OneFitsAll, and LLaTA take, on average, 28.2, 2.3, and 1.2 times longer than the modified models. Examples can be seen in Figure 3, where the green marks (ablation methods) are typically below the red one (LLM) and are positioned towards the left of the axis, indicating a lower computational costs and better forecasting performance. Other datasets and MSE metric refer to Figure 7 and Figure 8 in Appendix. In conclusion, the computational intensity of LLMs in time series forecasting tasks does not result in a corresponding performance improvement.
@@ 4.3 Does language model pretraining help performance on forecasting tasks? (RQ3)
Our evaluation in this section indicates that pretraining with language datasets is unnecessary for time series forecasting. To test whether the knowledge learned during pretraining meaningfully improves forecasting performance we experimented with different combinations of pretraining and finetuning LLaTA’s [19] language model on time series.
  • Pretrain + Finetune (Pre+FT). This is the original method, wherein a pretrained language model is finetuned on time series data. In the case of LLaTA, the base language model is frozen and low rank adapters (LoRA) are learned.
@@ 4.5 Do LLMs help with few-shot learning in forecasting? (RQ5)
While our results indicate that LLMs are not useful for time series forecasting, it is nonetheless possible that knowledge encoded in pretrained weights could help performance in few-shot settings where data are scarce. To evaluate whether this is the case we trained models and their ablations on 10% of each dataset. Specifically, we evaluated LLaMA in Time-LLM methods. The results for LLaMA, shown in Table 7, compared LLaMA with completely removing the LLM (w/o LLM). There was no difference, with each performing better in 8 cases. We conducted similar experiments with LLaTA, a GPT-2-based method. Our results in Table 8 indicate that our ablations can perform better than LLMs in few-shot scenarios.
@@
Table 8: In few-shot scenarios (10% dataset), Ablation methods perform much better than GPT-2 (LLaTA). Without LLMs, 12 out of 14 cases showed better performance.
@@ 4.6 Where does the performance come from? (RQ6)
The specific results, shown in Table 19 in the Appendix, indicate that a structure combining patching and attention, named “PAttn”, performs better than most other encoding methods on small datasets (with time stamps less than  1 million) and is even comparable to LLM methods. Its detailed structure, as shown in Figure 4, involves applying "instance norm" to the time series, followed by patching and projection. Then, one-layer attention enables feature learning between patches. For larger datasets, such as Traffic (~15 million) and Electricity (~8 million), a single-layer linear model with a basic transformer, named "LTrsf," performs better in encoding. In those methods, finally, time series embedding will be projected with a single linear layer to forecast. Details of other encoders is in Appendix subsection D.3.
@@ 5 Conclusion
In this paper we showed that despite the recent popularity of LLMs in time series forecasting they do not appear to meaningfully improve performance. We experimented with simple ablations, showing that they maintain or improve the performance of the LLM-based counterparts while requiring considerably less compute. Once more, our goal is not to suggest that LLMs have no place in time series analysis. To do so would likely prove to be a shortsighted claim. Rather, we suggest that the community should dedicate more focus to the exciting tasks could be unlocked by LLMs at the interface of time series and language such as time series reasoning [22], or social understanding [6].
@@ D.2 Baseline Hyper-Parameter Details
When reproducing the reference methods, we used the original repository’s hyper-parameters and model structures. In the ablation study, due to the smaller model parameters, we adjusted the learning rate or increased the batch size in some cases. All other training details remained identical with the reference methods. The specific training details and hyper-parameters are provided in the code and run scripts of the repository [^2] https://anonymous.4open.science/r/TS_Models-D1BC/. Note that the training process and hyper-parameters for the simple methods can also be accessed via this link.
D.3 Details of Simple Methods
To investigate the source of LLM method performance, we conducted further research on encoders. We used various encoders to encode time series data, followed by a linear layer to project the time series embeddings to the forecast. The encoder structure combining patching and attention is shown in Figure 4. In addition, we propose three different neural models (i) “LTrsf”: which performs better on larger datasets, directly projects the time series into a transformer for encoding; (ii) ‘̀D-LTrsf”; and (iii) “D-PAttn”: both of them decompose the time series into three sub-sequences and forecast each using the above two methods respectively, and then linearly combine the results for the final forecast. Across our results in Table 19, we note that even simpler models significantly outperform LLM-based time-series models. In detail, the LLM-based models, all combined we able to appear 33 times as the best and the second-best performer. However, ‘PAttn’ was outperforming them by appearing 34 times as the best and the second-best performer.

### v2

@@ Abstract
Large language models (LLMs) are being applied to time series forecasting. But are language models actually useful for time series? In a series of ablation studies on three recent and popular LLM-based time series forecasting methods, we find that removing the LLM component or replacing it with a basic attention layer does not degrade forecasting performance—in most cases, the results even improve! We also find that despite their significant computational cost, pretrained LLMs do no better than models trained from scratch, do not represent the sequential dependencies in time series, and do not assist in few-shot settings. Additionally, we explore time series encoders and find that patching and attention structures perform similarly to LLM-based forecasters. [^1] All resources needed to reproduce our work are available: https://github.com/BennyTMT/LLMsForTimeSeries
@@ 1 Introduction
Time series analysis is a critical problem across many domains, including disease propagation forecasting [8], retail sales analysis [3], healthcare [25, 16] and finance [30]. A great deal of recent work in time series analysis (constituting repositories with more than 1200 total stars on GitHub) has focused on adapting pretrained large language models (LLMs) to classify, forecast, and detect anomalies in time series [14, 49, 21, 4, 5, 31, 13, 43, 15]. These papers posit that language models, being advanced models for sequential dependencies in text, may generalize to the sequential dependencies in time series data. This hypothesis is unsurprising given language models are now pervasive in machine learning research. However, direct connections between language modeling and time series forecasting remain largely undefined. So to what extent is language modeling really beneficial for traditional time series tasks?
Our claim is simple but profound: popular LLM-based time series forecasters perform the same or worse than basic LLM-free ablations, yet require orders of magnitude more compute. Derived from extensive ablations, this reveals a worrying trend in contemporary time series forecasting literature. Our goal is not to imply that language models will never be useful for time series. In fact, recent works point to many exciting and promising ways that language and time series interact, like time series reasoning [24, 7, 44, 41, 36], social understanding [6] and financial reasoning [35, 19]. Rather, we aim to highlight surprising findings that existing methods do very little to use the innate reasoning power of pretrained language models on established time series tasks.
@@
  • We study the impact of an LLM’s pretraining by re-initializing their weights prior to forecasting. We find that this has no impact on forecasting performance. Additionally, in shuffling input time series, we find no evidence the LLMs successfully transfer sequence modeling abilities from text to time series and no indication they help in few-shot settings.
@@ 2 Related Work
Here, we summarize the key works relevant to LLM-based time series models. They can be broadly classified into three sections: (i) time series forecasting using LLMs; (ii) encoders in LLM time series models; and (iii) smaller and efficient neural models for time-series.
Time Series Forecasting Using LLMs. Recently, with the development of Large Language Models (LLMs) [10, 28, 33] and their demonstrated multi-modal capabilities, more researchers have successfully applied LLMs to time series forecasting tasks [13, 15, 5, 4]. Chang et al.,  [5] used finetuning the transformer module and positional encoding in GPT-2 to align pre-trained LLMs with time series data for forecasting tasks. Zhou et al. [49] proposed a similar finetuning method, named “OneFitAll”, for time series forecasting with GPT-2. Additionally, Jin et al. [14] introduced a reprogramming method to align LLM’s Word Embedding with time series embeddings, showing good representation of time series data on LLaMA [33]. Similarly, CALF [21] and Test [31] adapted word embeddings to enable LLMs to forecast time series data effectively. In addition to time-series forecasting models, Liu et al. [22] show that these models can be extended to classifying health-time series, such as heart-rate and daily-footsteps. These models have also been shown to outperform supervised neural models in few-shot settings.
Encoders in LLM Time Series Models. In order for an LLM to learn from text it must first be discretized and encoded as word tokens which are $1\times d$ vectors [10, 28, 33]. Similarly, LLM-based methods for time series learn discrete time series tokens. One method is to segment the time series into overlapping patches, which effectively shortens the time series while retaining its features [14, 49, 5, 4, 27, 26]. Another method involves decomposing a method based on its trend, seasonal components, and residual components [4, 27]. Lastly, Liu et al. [21] feed the multivariate time series using a Transformer to enable different channels to learn the dynamics of other channels. These embedding procedures are followed by a linear neural network layer that projects the time series encoding to the same dimensions used by the pre-trained LLM.
Small and Efficient Neural Forecasters. In addition to LLMs, there has been a large body of research on smaller yet efficient frameworks that outperform their bulky counterparts in time series forecasting [18, 46, 32, 23, 2]. For example, Zeng et al. [45] present DLinear, an incredibly simple model that combines decomposition techniques and achieves better forecasting performance than state-of-the-art transformer-based time series architectures at the time, such as Informer [47], FEDformer [48], and Autoformer [39]. Furthermore, Xu et al. [42] introduces a lightweight model with only 10k parameters, which captures both amplitude and phase information in the time-series to outperform transformer-based models.
@@ 3 Experimental Setup
CALF [21] GPT-2 LoRA Fine-Tune
@@ 3.1 Reference Methods for Language Models and Time Series
  • CALF [21]: CALF embeds the input time series by treating each channel as a token. One half of the architecture is a “textual branch” which uses cross attention to align the time series representation with a low dimensional representation of the language model’s word embeddings. This representation is then passed through a pretrained, frozen language model to obtain a “textual prediction”. Simultaneously, a “temporal” branch learns a low-rank adapter for a pretrained language model based on the input time series to produce a “temporal prediction” which is used for inference. The model includes additional loss terms that enforce similarity between these representations.
@@ 4 Results
CALF [21] Model $\rightarrow$ CALF w/o LLM LLM2Attn LLM2Trsf From Original Paper
@@
Table 3: Forecasting performance of all models – Time-LLM, CALF, and OneFitsAll and results from our ablations. All results are averaged across different prediction lengths, though full results are available in Appendix E.1. Results in Red denote the best-performing model. # Wins refers to the number of times the method performed best, and # Params is the number of model parameters. “-” means the dataset is not included in the original paper.
@@ 4.1 Do pretrained language models contribute to forecasting performance? (RQ1)
Our results show that pretrained LLMs are not useful for time series forecasting tasks yet. Overall, as shown in Table 3, across 13 datasets and two metrics, ablations out perform Time-LLM methods in $/26$ 26 cases, CALF in $22/26$ cases, and OneFitsAll in $19/26$ cases. We averaged results over different predicting lengths, as in [49, 14, 21]. Across all prediction lengths (thirteen datasets and four prediction lengths) ablations outperformed Time-LLM, CALF, and OneFitsAll in $35/40$, $31/40$, and $29/40$ cases as measured by MAE, respectively. To ensure a fair comparison, we also report results from each method’s original paper alongside our replication. For specific results refer to Appendix E.1. To better evaluate the effectiveness of LLMs and ablation methods, we include 95% bootstrapped confidence intervals for each task. In tasks where LLMs performed better, such as OneFitsAll with ETTh1, shown in  Figure 2, there is still substantial overlap in the confidence intervals with the ablation method “w/o LLM” in MAE. Other datasets results and MSE metrics are shown in Figure 5 and Figure 6 in the Appendix. To summarize, our results on the evaluation above, it is hard to conclude that LLMs are effective in time series forecasting.
Figure 2: In the above examples, only OneFitsAll “w/ LLM” performs better than the ablation methods on ETTh1, but there is substantial overlap in bootstraped confidence intervals. The figures show the comparison of OneFitsAll, CALF, and Time-LLM using LLMs and ablations (i.e., w/o LLM, LLM2Attn, and LLM2Trsf) on ETTh1, ETTm2, and Electricity, and the vertical dashed lines represent the results from the original work. Others Figures for MSE and other datasets are available in Figure 5 and Figure 6 in the Appendix.
Figure 3: Ablation methods consume less time for inference while providing better forecasting performance. The figure above shows the inference time and prediction accuracy of Time-LLM, OneFitsAll, and CALF on ETTm2, Traffic, and Electricity datasets, averaged across prediction lengths. For more datasets and MSE metrics refer to Figure 7 and Figure 8 in the Appendix.
Method Time-LLM (LLaMA) OneFitsAll (GPT-2) CALF (GPT-2)
@@ 4.2 Are LLM-based methods worth the computational cost? (RQ2)
In the previous section, we showed that LLMs do not meaningfully improve performance on time series forecasting tasks. Here, we evaluate the computational intensity of these methods with their nominal performance in mind. The language models in our reference methods use hundreds of millions and sometimes billions of parameters to perform time series forecasting. Even when the parameters of the language models are frozen they still contribute to substantial overhead during training and inference. For instance, Time-LLM has 6642 M parameters and takes 3003 minutes to train on the Weather dataset whereas ablation methods have only 0.245 M parameters and take 2.17 minutes on average. Information about training other methods on ETTh1 and Weather datasets are shown in Table 4. In the case of inference time, we divide by the maximum batch size to give an estimate of inference time per example. Time-LLM, OneFitsAll, and CALF take, on average, 28.2, 2.3, and 1.2 times longer than the modified models. Examples can be seen in Figure 3, where the green marks (ablation methods) are typically below the red one (LLM) and are positioned towards the left of the axis, indicating a lower computational costs and better forecasting performance. Other datasets and MSE metric refer to Figure 7 and Figure 8 in Appendix. In conclusion, the computational intensity of LLMs in time series forecasting tasks does not result in a corresponding performance improvement.
@@ 4.3 Does language model pretraining help performance on forecasting tasks? (RQ3)
Our evaluation in this section indicates that pretraining with language datasets is unnecessary for time series forecasting. To test whether the knowledge learned during pretraining meaningfully improves forecasting performance we experimented with different combinations of pretraining and finetuning CALF’s [21] language model on time series.
  • Pretrain + Finetune (Pre+FT). This is the original method, wherein a pretrained language model is finetuned on time series data. In the case of CALF, the base language model is frozen and low rank adapters (LoRA) are learned.
@@ 4.5 Do LLMs help with few-shot learning in forecasting? (RQ5)
While our results indicate that LLMs are not useful for time series forecasting, it is nonetheless possible that knowledge encoded in pretrained weights could help performance in few-shot settings where data are scarce. To evaluate whether this is the case we trained models and their ablations on 10% of each dataset. Specifically, we evaluated LLaMA in Time-LLM methods. The results for LLaMA, shown in Table 7, compared LLaMA with completely removing the LLM (w/o LLM). There was no difference, with each performing better in 8 cases. We conducted similar experiments with CALF, a GPT-2-based method. Our results in Table 8 indicate that our ablations can perform better than LLMs in few-shot scenarios.
@@
Table 8: In few-shot scenarios (10% dataset), Ablation methods perform much better than GPT-2 (CALF). Without LLMs, 12 out of 14 cases showed better performance.
@@ 4.6 Where does the performance come from? (RQ6)
The specific results, shown in Table 18 in the Appendix, indicate that a structure combining patching and attention, named “PAttn”, performs better than most other encoding methods on small datasets (with time stamps less than  1 million) and is even comparable to LLM methods. Its detailed structure, as shown in Figure 4, involves applying "instance norm" to the time series, followed by patching and projection. Then, one-layer attention enables feature learning between patches. For larger datasets, such as Traffic (~15 million) and Electricity (~8 million), a model named "LTrsf," using the encoder from CALF [21], performs better. In those methods, finally, time series embedding will be projected with a single linear layer to forecast. Details of other encoders is in Appendix subsection D.3.
@@ 5 Conclusion
In this paper we showed that despite the recent popularity of LLMs in time series forecasting they do not appear to meaningfully improve performance. We experimented with simple ablations, showing that they maintain or improve the performance of the LLM-based counterparts while requiring considerably less compute. Once more, our goal is not to suggest that LLMs have no place in time series analysis. To do so would likely prove to be a shortsighted claim. Rather, we suggest that the community should dedicate more focus to the exciting tasks could be unlocked by LLMs at the interface of time series and language such as time series reasoning [24, 7, 44, 36], or social understanding [6].
6 Acknowledgements
We thank the University of Virginia’s Research Computing team for maintaining the excellent high-performance computing resources that allowed us to conduct this research. This research was supported in part by NSF IIS-1901386, NSF CAREER IIS-2142794, NIH R01MH125179, Bill & Melinda Gates Foundation (INV-004841), and the Office of Naval Research (#N00014-21-1-2154).
@@ D.2 Baseline Hyper-Parameter Details
When reproducing the reference methods, we used the original repository’s hyper-parameters and model structures. In the ablation study, due to the smaller model parameters, we adjusted the learning rate or increased the batch size in some cases. All other training details remained identical with the reference methods. The specific training details and hyper-parameters are provided in the code and run scripts of the repository [^2] https://github.com/BennyTMT/LLMsForTimeSeries. Note that the training process and hyper-parameters for the simple methods can also be accessed via this link.
D.3 Details of Encoder Exploration and Simple Methods
To investigate the source of LLM method performance, we conducted further research on encoders. We used various encoders to encode time series data, followed by a linear layer to project the time series embeddings to the forecast. The encoder structure combining patching and attention is shown in Figure 4. The key difference between PAttn and PatchTST [26] is the absence of position embedding and the Feed Forward in PatchTST [26], or, more specifically, replacing the Transformer Encoder with a simple single-layer Attention structure. It could be a simple yet effective method to help evaluate the trade-off between cost and performance for newly proposed methods.
In addition, we propose three different neural models (i) “LTrsf”: which performs better on larger datasets, uses CALF ’s [21] encoder without cross-modal attention. This could be a potential source of CALF ’s performance; (ii) “D-LTrsf”; and (iii) “D-PAttn”: both of them decompose the time series into three sub-sequences and forecast each using the above two methods respectively, and then linearly combine the results for the final forecast. Across our results in Table 18, we note that even simpler models significantly outperform LLM-based time-series models. In detail, the LLM-based models, all combined we able to appear 33 times as the best and the second-best performer. However, “PAttn” was outperforming them by appearing 34 times as the best and the second-best performer. In addition, the comparison between “PAttn”, “LTrsf” and other state-of-the-art methods is shown in Table 19.